Question: Jacob has $6.00 in dimes and quarters. He has ten more quarters than dimes. How many dimes and quarters does he have?


I don't know how to set up these kinds of problems... I tried this, but I don't really know what to do after this---- $6.00= (10+q)dx ----- It's probably wrong but can you please explain how you set it up and why?? thank you SOOO much!!!!

Answer by ankor@dixie-net.com(22740)

Do it this way
Let d = no. dimes
Let q = no. of quarters
:
Write an equation for each statement
:
"Jacob has $6.00 in dimes and quarters."
.10d + .25q = 6.00
:
" He has ten more quarters than dimes."
q = d + 10
:
In the first equation. replace q with (d+10)
.10d + .25(d+10) = 6
.10d + .25d + 2.5 = 6
.35d = 6 - 2.5
.35d = 3.5
d = 3.5/.35
d = 10 dime
then obviously we have 10 more quarters or 20 quarters
:
:
Find the total using these value: .10(10) + .25(20) =
:
:
This is a method you can use on all these type problems